OpenStudy (anonymous):

Help with Expressions?? Using complete sentences, describe how you would check the product of two rational expressions. Create and demonstrate your own example of this concept. Means a LOT! THANK YOU!!

OpenStudy (nincompoop):

check the product of two rational expressions hmm

OpenStudy (nincompoop):

\[\frac{a}{b}\times \frac{c}{d}=\frac{ab}{cd}\]

OpenStudy (anonymous):